OpenStudy (anonymous):

Fractions: 16 2/9 - 2 2/3

myininaya (myininaya):

\[16 \frac{2}{9}-2\frac{2}{3}=\frac{9(16)+2}{9}-\frac{3(2)+2}{3}\] Write both as improper fractions

OpenStudy (anonymous):

hahaha answer its due to marrow

myininaya (myininaya):

try the simplify the numerators of both fractions and then we will go from there

OpenStudy (anonymous):

13 5/9

myininaya (myininaya):

\[\frac{144+2}{9}-\frac{6+2}{3}\] Notice I performed that multiplication first. \[\frac{146}{9}-\frac{8}{3}\] Now we need to find a common denominator both 9 and 3 go into 9 so we will write both fractions such that their denominators are both 9 \[\frac{146}{9}-\frac{8}{3} \cdot \frac{3}{3}\] \[\frac{146}{9}-\frac{24}{9}\] Combine the fractions since they have the same denominator now \[\frac{146-24}{9}\] \[\frac{122}{9}\] 13 ---- 9| 122 -9 ------ 32 -27 ---- 5 yes jordon that is right :)

OpenStudy (anonymous):

yay